# Sample Compression for Self-Certified Continual Learning

### v3

@@ 4 Experiments
Datasets and continual learning settings.  We carry-out experiments on MNIST lecun2010mnist , Fashion-MNIST fashionmnist , EMNIST (characters) cohen2017emnistextensionmnisthandwritten , PermutedMNIST goodfellow2013empirical , Rotated MNIST, CIFAR10 cifar10cifar100 , CIFAR100 cifar10cifar100 and TinyImageNet tiny-imagenet . We report results that include Class-Incremental (CI), Task-Incremental (TI), and Domain-Incremental (DI) continual learning settings. The dataset stats regarding how many classes/data items are available per task is given in Appendix H.
@@
Baselines.  To assess the effectiveness of our proposed method CoP2L, we benchmark against several strong and widely adopted baselines across Class-Incremental, Task-Incremental, and Domain-Incremental continual learning settings. For each continual learning set-up we include baselines that yield the best performance in each category according to the Avalanche benchmark lomonaco2021avalanche .
For the Class-Incremental setting, we use the baselines Replay rolnick2019experience , iCaRL rebuffi2017icarl , and GDumb prabhu2020gdumb . Replay in particular is considered a strong baseline DellaLibera2023CLMASRAC due to its simplicity and effectiveness. In the Task-Incremental setting, we compare against Replay rolnick2019experience , LaMAML gupta2020look and Learning without Forgetting (LwF) li2017learning . These methods are well suited for TI scenarios where task identity is known at test time. In the Domain-Incremental setting, we evaluate against Replay rolnick2019experience , Less Forgetful Learning (LFL) jung2018less , Elastic Weight Consolidation (EWC) kirkpatrick2017overcoming , and Synaptic Intelligence (SI) zenke2017continual .
Implementation.  All our experiments are conducted using the Avalanche continual learning framework avalanchelib , which provides a unified interface for datasets, benchmarks, and baseline implementations. We also use the official ContinualAI benchmark repository [^1] https://github.com/ContinualAI/continual-learning-baselines for consistent implementation and fair comparisons. The code for the experimental results is available in the supplemental material.
Model architecture and training details.  We use the SimpleMLP architecture and customized SimpleCNN architectures (SimpleMLP and SimpleCNN from the Avalanche toolkit) for the MNIST, FMNIST, and EMNIST experiments. All models are trained using the SGD optimizer. The details of the models architectures and the hyperparameters used can be found in Appendix G. For experiments on more complex datasets such as CIFAR10, CIFAR100, and TinyImageNet, we use the pretrained feature extractor Dino caron2021emerging using ViT dosovitskiy2020image and ResNet50 he2016deep backbones, followed by a trainable MLP head. This choice is motivated by the insight that CoP2L can further benefit from stronger feature representations, especially in settings with limited data per class or high dataset complexity. By leveraging Dino’s obtained embeddings, we enhance the quality of the input representations, which we empirically observe to yield good performance. Results with standard deviations across seeds and total training times are reported in Appendix F. We used a single NVIDIA 4090 GPU for our experiments.
Accuracy and Forgetting.  We denote the accuracy obtained on a task $t$, after training the model for $T$ tasks with $\text{Accuracy}_{T,t}$. The accuracy we report is defined as,
. $\textit{Average Accuracy}_{T}=\frac{1}{T}\sum_{t=1}^{T}\text{Accuracy}_{T,t}$, where $t$ is for the task index. .
We also report the average forgetting at task $T$ which is defined as
$\textit{Average Forgetting}_{T}=\frac{1}{(T-1)}\sum_{t=1}^{T-1}[\text{Accuracy}_{t,t}-\text{Accuracy}_{T,t}],$ which is calculated when $T>1$.
4.1 Main Results
In Table 1, we present the average accuracy and average forgetting values defined in the section above, obtained on MNIST, FMNIST and EMNIST dataset, both with an MLP and a CNN architecture, where the networks are trained from scratch. In Appendix F, we also provide the standard deviations as well. We observe that with the MLP architecture CoP2L is able to outperform the baselines. Furthermore, we also observe that CoP2L is able to obtain competitive results when employed with a CNN architecture as well. In Appendix E, we also provide an analysis of the performance of CoP2L for varying buffer sizes.
In Table 2, we extended our results to cover CIFAR10, CIFAR100, and TinyImageNet datasets, for both Class-Incremental and Task-Incremental settings. We also provide results for Domain-Incremental setting with Permuted-MNIST goodfellow2013empirical and Rotated-MNIST datasets.
Table 1: Summary of average accuracies and forgetting values obtained on MNIST, FMNIST, EMNIST (5 tasks, Class-Incremental), and on Permuted-MNIST and Rotated-MNIST (10 tasks, Domain-Incremental).
Method / Dataset MNIST FMNIST EMNIST
Acc / For Acc / For Acc / For
CoP2L (MLP) 95.1 / 3.4 84.2 / 12.1 78.0 / 16.7
Replay (MLP) 94.5 / 6.1 84.0 / 15.9 77.2 / 22.1
GDumb (MLP) 91.7 / 3.8 81.8 / 10.4 62.6 / 16.0
CoP2L (CNN) 96.6 / 2.2 87.4 / 11.1 87.9 / 8.4
Replay (CNN) 96.9 / 3.3 87.6 / 13.3 87.7 / 12.3
GDumb (CNN) 94.4 / 3.0 85.0 / 8.8 85.1 / 7.5
Method / Dataset PermMNIST RotMNIST
Acc / For Acc / For
CoP2L (MLP) 92.3 / 6.0 88.3 / 9.8
Replay (MLP) 91.5 / 6.4 88.2 / 10.0
LFL (MLP) 83.5 / 9.3 51.6 / 45.7
EWC (MLP) 83.7 / 5.2 56.5 / 35.4
SI (MLP) 86.7 / 10.7 57.7 / 43.9
@@ 4.1 Main Results
Table 2: Summary of average accuracies and forgetting values obtained on various continual learning tasks on CIFAR10, CIFAR100 and TinyImageNet datasets. CI denotes Class-Incremental, TI denotes Task-Incremental continual learning settings.
CIFAR10 (5 tasks), CI CIFAR100 (10 tasks), CI CIFAR100 (20 tasks), CI TinyImNet (40 tasks), CI
@@
CoP2L-Dino (ViT) 94.45 3.36 73.83 16.76 72.50 16.96 51.13 36.81
Replay-Dino (ViT) 94.01 6.01 68.51 30.44 69.23 29.60 47.66 48.15
iCaRL-Dino (ViT) 93.24 1.55 61.50 14.53 64.11 15.25 20.78 52.95
GDumb-Dino (ViT) 94.15 3.39 72.73 11.70 72.05 11.38 54.53 14.79
CoP2L-Dino (RN50) 81.13 5.55 49.18 21.99 49.98 24.32 32.74 36.40
Replay-Dino (RN50) 81.83 17.09 49.40 42.07 49.91 44.41 28.43 63.26
iCaRL-Dino (RN50) 74.56 1.54 45.97 8.71 46.87 6.32 33.13 10.95
GDumb-Dino (RN50) 80.70 8.95 43.83 20.31 43.78 19.03 24.04 21.82
CIFAR10 (5 tasks), TI CIFAR100 (10 tasks), TI CIFAR100 (20 tasks), TI TinyImNet (40 tasks), TI
CoP2L-Dino (ViT) 99.01 -0.03 95.15 0.92 97.10 0.71 94.28 1.22
Replay-Dino (ViT) 99.24 -0.03 96.59 0.23 98.06 0.41 95.82 0.36
LaMAML-Dino (ViT) 99.18 0.06 95.77 0.40 97.70 0.33 94.90 1.15
LwF-Dino (ViT) 97.53 2.08 95.26 1.06 95.73 2.15 85.79 9.09
CoP2L-Dino (RN50) 96.17 0.79 86.58 2.26 90.88 2.37 88.75 2.13
Replay-Dino (RN50) 96.96 0.42 87.59 0.64 93.13 0.16 90.53 0.74
LaMAML-Dino (RN50) 96.71 0.66 83.13 6.81 90.57 3.35 87.25 4.55
LwF-Dino (RN50) 95.90 0.15 79.29 0.67 87.94 0.09 85.15 0.90
Figure 1: Example demonstration of emprical upper bounds on test errors, on Class-Incremental (top-left CIFAR10, top-right CIFAR100) and Task-Incremental (bottom-left CIFAR10, bottom-right CIFAR100) datasets.
@@
Empirical study of the CoP2L generalization bound.  In Figure 1, we showcase the empirical upper bounds on test errors. The details on the computation of the bound can be found in the Appendix  A. We observe that the bound values are non-trivial and generally follow similar trends as the actual error we report on the test set, while the estimate for the bound is calculated exclusively on the training set. In Figure 1, we provide plots for the bounds for the models trained with a buffer size of 2000, for Class-Incremental and Task-Incremental settings. CoP2L achieves particularly tight generalization bounds in the Task-Incremental setting, where forgetting is inherently reduced due to access to task identities and the use of a separate classifier head for each task. These conditions allow the model to better compress task-specific knowledge, leading to both stronger theoretical guarantees and improved empirical performance. We also observe that with the ViT backbone, the bounds are significantly tighter than those obtained with the ResNet50 backbone. This is likely because the ViT backbone provides more structured and information-rich representations, which are easier to compress effectively, thereby enhancing the performance of CoP2L.
We also give the results for all buffer sizes for the three MNIST datasets, and also bound estimates for the CIFAR10, CIFAR100, and ImageNet datasets (TinyImageNet tiny-imagenet , ImageNette imagenette and ImageWoof Howard_Imagewoof_2019 ) as well as bounds for the Domain-Incremental setting in Appendix D.
4.2 Analysis of CoP2L
In this section we provide an analysis of CoP2L with an ablation study on loss terms, a sensitivity analysis with respect to task ordering, and buffer size. Due to space limitations the figures are provided in the appendix, however the main results are mentioned here in the text.
Ablation on CoP2L loss terms  We have conducted an ablation study to understand the effect of additional weighting that counteracts class imbalance that arises in the continual learning setup when new tasks are introduced (Standard replay counteracts this by extending the batch size to have equal class representation). In the modified P2L (Algorithm 2), we incorporate an additional weighting term over classes to adapt P2L for class-balanced continual learning. In order to assess whether this idea is indeed helpful, we have run an experiment on the MNIST dataset with and without the additional weighting idea. In Appendix E, in Figure 10, we show this experiment. For both accuracy (left panel) and forgetting (right panel), the inclusion of the additional loss weighting term significantly improves the results.
@@ 4.2 Analysis of CoP2L
Analysis on buffer size.  In Appendix E, in Figure 12, we compare the accuracy and forgetting obtained with different replay buffer sizes of (1000, 2000, 3000, 4000, 5000) samples for CoP2L and replay. This represents roughly 8% to 40% of the per-task dataset size, for MNIST and Fashion-MNIST, and 10% to 52% for EMNIST. For each replay buffer size, we report the variation of accuracy obtained at the end of training averaged over different seeds, and test accuracy for different tasks.
For the MNIST dataset, on average CoP2L significantly outperforms replay (as evidenced by the variation over seeds). For the Fashion-MNIST and EMNIST datasets, when using small buffer sizes, the average accuracy obtained with our CoP2L is significantly better. For larger buffer sizes, in general CoP2L performs comparably or slightly better. We also note for all three datasets, CoP2L significantly outperforms replay in terms of forgetting. We attribute this due to the fact that CoP2L is able to intelligently select examples from the replay buffer. Namely, CoP2L is able to focus on examples that are most representative of the earlier tasks, which we observe to help significantly in mitigating the deterioration over earlier tasks as evidenced by the results presented in the right column in Figure 12.
5 Limitations and Broader Impact Statement
We observe with our experiments that the performance of CoP2L is dependent to the ability of the feature extractor (backbone). In case a weak backbone is employed, CoP2L could necessitate a larger compression set (e.g. more data points per task). We have also seen that in those cases CoP2L might take longer to converge. We noted in our experiments that using deep pretrained networks such as ViT results in good performance, by increasing the compression ability. More broadly, we do not believe there is any particular negative social impact to note for our proposed methodology. On the contrary, we believe that our algorithm could be instrumental for improving trustworthiness of continual learning / lifelong learning in practice.
6 Conclusions
In this paper, we have proposed CoP2L, an algorithm rooted in the sample compression theory for self-certified continual learning. To the best of our knowledge, this is the first attempt in the literature to employ sample compression theory within the continual learning context. We provided sample compression bounds for CoP2L and verified empirically that they are non-vacuous and informative. We furthermore showed that on several challenging datasets, that our approach yields comparable or better results compared several strong baselines, while also providing learning guarantees, for trustworthy continual learning. As the combination of Pick-to-Learn with experience replay leads to a successful learning scheme, we foresee that combining Pick-to-Learn with other continual learning approaches could also lead to new compelling ways of obtaining self-certified predictors.
Acknowledgements
Mathieu Bazinet is supported by a FRQNT B2X scholarship (343192, doi: https://doi.org/10.69777/343192). Pascal Germain, and Cem Subakan are supported by Discovery Grants RGPIN-2020-07223 and RGPIN 2023-05759, respectively. This research was also enabled in part by support provided by Calcul Québec (https://www.calculquebec.ca/) and the Digital Research Alliance of Canada (https://alliancecan.ca/en).
Appendix A Compression and Reconstruction Functions
In line with the prevalent literature, we introduced the sample compression framework (Section 2) based on a compression function providing a compression set and optionally a message. In the case of CoP2L, the reconstruction scheme relies on two compression sets $(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}})$ and a message pair $(\mu_{1},\mu_{2})$, which we explain later in this section. [^2] The idea of using multiple compression sets appeared in [40] and [39], but in a setting without a message.

### v4

@@ 4 EXPERIMENTS
In this section, we present experiments with CoP2L to demonstrate the tightness of our new generalization bounds for continual learning in two different settings. Afterward, we compare CoP2L to established baseline approaches to show that although our algorithm was first designed to obtain tight generalization bounds, its performances are still very competitive with established continual learning schemes.
Datasets and continual learning settings.
We carry out experiments on CIFAR10 (Krizhevsky, 2012), CIFAR100 (Krizhevsky, 2012) and TinyImageNet (mnmoustafa and Ali, 2017). We report results in Class-Incremental (CI) and Task-Incremental (TI) continual learning settings, as defined in the Avalanche framework (Carta et al., 2023).
Baselines.
@@
To assess the effectiveness of our proposed method CoP2L, we benchmark against several strong and widely adopted baselines across Class-Incremental and Task-Incremental learning settings.
For the Class-Incremental setting, we use the baselines Replay (Rolnick et al., 2019), Dark Experience Replay (DER) (Buzzega et al., 2020), iCaRL (Rebuffi et al., 2017b), GDumb (Prabhu et al., 2020), Contrastive Continual Learning via Importance Sampling (CCLIS) (Li et al., 2024) and Coreset Selection via Reducible Loss (CSReL) (Tong et al., 2025)
In the Task-Incremental setting, we report Replay, DER, LaMAML (Gupta et al., 2020), Learning without Forgetting (LwF) (Li and Hoiem, 2017), CCLIS and CSReL.
(a) Class-incremental setting
(b) Task-incremental setting
Figure 2: Illustration of the behavior of the bound on CIFAR10 with 5 tasks.
(a) Class-incremental setting
(b) Task-incremental setting
Figure 3: Illustration of the behavior of the bound on CIFAR100 with 20 tasks.
@@
Table 1: Summary of average accuracies and forgetting values obtained on various continual learning tasks on CIFAR10, CIFAR100 and TinyImageNet datasets.
(a) Class-Incremental experiments
CIFAR10 (5 tasks) CIFAR100 (10 tasks) CIFAR100 (20 tasks) TinyImNet (40 tasks)
@@
CoP2L (ViT) 94.45 2.10 73.49 17.06 70.56 21.15 51.59 36.32
Finetuning (ViT) 27.55 89.74 12.87 93.27 8.49 94.64 2.83 95.36
Replay (ViT) 94.00 6.11 68.71 30.02 69.21 29.52 47.49 48.39
DER (ViT) 95.03 3.79 72.29 25.11 77.03 19.27 53.92 41.77
iCaRL (ViT) 88.60 3.13 65.54 13.64 62.36 15.88 21.06 34.49
GDumb (ViT) 94.16 3.16 72.51 11.53 72.35 11.77 53.93 15.11
CCLIS (ViT) 93.15 4.44 67.95 23.68 66.81 24.42 48.55 32.26
CoP2L (RN50) 80.98 5.84 49.53 22.31 46.74 22.54 33.30 35.93
Finetuning (RN50) 19.63 97.22 9.36 89.93 4.88 92.81 1.80 78.66
Replay (RN50) 81.38 17.72 49.41 42.12 49.44 44.96 28.38 63.36
DER (RN50) 82.35 15.31 49.37 44.01 57.53 33.37 28.77 64.17
iCaRL (RN50) 74.49 1.79 46.06 8.50 46.47 6.74 33.89 10.42
GDumb (RN50) 80.84 9.35 43.37 20.23 43.60 19.93 24.29 21.29
CCLIS (RN50) 76.89 12.32 44.22 23.23 39.10 25.78 18.32 40.40
CSReL (RN50) 34.50 26.44 38.75 42.12 32.57 40.23 12.84 73.69
(b) Task-Incremental experiments
CIFAR10 (5 tasks) CIFAR100 (10 tasks) CIFAR100 (20 tasks) TinyImNet (40 tasks)
CoP2L (ViT) 99.04 -0.00 95.03 0.90 97.18 0.56 93.81 1.50
Finetuning (ViT) 99.15 0.23 95.17 1.84 97.57 0.86 93.62 2.61
Replay (ViT) 99.29 -0.01 96.43 0.34 98.09 0.25 95.66 0.50
DER (ViT) 99.15 0.21 96.22 0.71 98.08 0.31 95.08 1.13
LaMAML (ViT) 99.25 0.05 95.58 0.58 97.84 0.26 94.82 0.79
LwF (ViT) 97.89 1.69 95.09 0.98 95.74 2.15 85.18 9.72
CCLIS (ViT) 98.81 0.48 90.99 4.49 94.86 2.88 86.82 6.14
CoP2L (RN50) 96.22 0.56 86.78 2.01 90.89 2.27 88.57 2.46
Finetuning (RN50) 96.60 1.00 88.23 1.78 93.25 1.27 90.62 1.38
Replay (RN50) 96.88 0.42 87.61 0.86 92.89 0.47 90.03 1.26
DER (RN50) 96.57 1.09 88.29 1.72 93.43 1.11 90.66 1.12
LaMAML (RN50) 96.65 0.73 83.57 6.18 89.89 4.12 87.20 4.53
LwF (RN50) 95.80 0.33 79.39 0.47 87.73 0.33 85.03 0.89
CCLIS (RN50) 95.45 0.79 78.55 6.76 83.51 6.60 75.65 14.12
CSReL (RN50) 65.46 13.66 69.14 15.17 63.52 14.49 50.22 22.34
All the details about the experiments can be found in Appendix˜E, alongside supplementary Class Incremental experiments on MNIST (LeCun et al., 2010), FashionMNIST (FMNIST) (Xiao et al., 2017) and EMNIST (Cohen et al., 2017). Moreover, we present Domain-Incremental (DI) experiments on PermutedMNIST (Goodfellow et al., 2014) and RotatedMNIST (Ben-David et al., 2010). We report all the results in Appendix E.3, including the standard deviations and the training times for each method.
Accuracy and Forgetting.
We denote $A(t,\theta_{T})$ the accuracy obtained on a task $t$ of a model $f_{\theta_{T}}$ trained on $T$ tasks. We report the average accuracy over $T$ tasks,
$\overline{A}(\theta_{T})\,{=}\,\frac{1}{T}\sum_{t=1}^{T}A(t,\theta_{T})\,$,
and the average forgetting at task $T$,
$\overline{F}(T)=\frac{1}{T-1}\sum_{t=1}^{T-1}[A(t,\theta_{t})-A(t,\theta_{T})]$}\,.
4.1 Study of our generalization bound
In Figures 2 and 3, we present our generalization bound on the true risk for each task for CIFAR10 with 5 tasks and CIFAR100 with 20 tasks. These bounds are presented for Vision Transformer (ViT) (Dosovitskiy et al., 2021) networks and ResNet50 (RN50) (He et al., 2016) networks trained with a buffer size of 2000 under both Class-Incremental and Task-Incremental settings. We present the details for the computation of the bound in Appendix˜B.
@@ 4.1 Study of our generalization bound
We observe that although the bounds are calculated exclusively on the training set, their values are non-vacuous and follow the test set error trends. CoP2L achieves particularly tight generalization bounds in the Task-Incremental setting, where forgetting is inherently reduced due to access to task identities and the use of a separate classifier head for each task. These conditions allow the model to better compress task-specific knowledge, leading to both stronger theoretical guarantees and improved empirical performance. We also observe that with the ViT backbone, the bounds are significantly tighter than those obtained with the ResNet50 backbone. This is likely because the ViT backbone provides more structured and information-rich representations, which are easier to compress effectively, thereby enhancing the performance of CoP2L. Furthermore, the tightness of the bounds improves when the size of the dataset increases, which explains that the bounds on CIFAR10 (5000 samples per class) are noticeably tighter than the bounds on CIFAR100 (about 500 samples per class).
In addition, the complexity of the incremental scenario plays a role: CIFAR10 with 5 tasks involve learning two classes at a time, whereas CIFAR100 with 10 tasks require learning ten classes simultaneously, further contributing to looser bounds. Generalization bounds for multiple combinations of datasets and architectures are reported in Section˜E.4.
4.2 Study of the algorithm’s performance
Table˜1 presents results on CIFAR10, CIFAR100 and TinyImageNet datasets, for both Class-Incremental and Task-Incremental settings. We observe that in Class-Incremental settings, CoP2L is competitive with most baselines in terms of accuracy while maintaining low forgetting. Alongside these baselines, we present results for finetuning a model without a replay buffer, which evidently leads to a drastic drop in performance. In Task-Incremental settings, our method remains competitive. We note that even simple finetuning performs well in this setting. This behavior is expected, since adding a new classification head for each task naturally reduces forgetting. All the while being a competitive continual learning approach, CoP2L offers the additional advantage of certifying the generalization loss of the learned predictor. Further experiments on Imagenette and Imagewoof (Howard, 2019a) are presented in Appendix E. We do not present results for the ViT with CSReL, as the method is designed for small datasets ($\approx$1000 datapoints), and becomes prohibitively more expensive when the dataset and the model grow in size. With the ResNet50, running CSReL was 5 to 32 times longer than executing CoP2L, while achieving much worse accuracy.
Finally, we refer the reader to Appendix F for an extended analysis of CoP2L behavior, including an ablation study on the loss terms, a sensitivity analysis with respect to task ordering, an analysis of the impact of buffer size, an ablation study on the use of early stopping with the bound and an evaluation of memory cost. Furthermore, we report experiments that CoP2L is able to mitigate forgetting more than the other methods while still retaining plasticity, and is able to achieve a balanced performance over past tasks.
5 CONCLUSION
We proposed CoP2L, an algorithm rooted in the sample compression theory, for self-certified continual learning. To the best of our knowledge, this is the first attempt to employ sample compression theory within the continual learning context. We provided sample compression bounds for CoP2L and verified empirically that they are non-vacuous and informative. We furthermore showed that on several challenging datasets, our approach yields comparable results when compared to several strong baselines, while also providing learning guarantees for trustworthy continual learning. As the combination of Pick-to-Learn with experience replay leads to a successful learning scheme, we foresee that combining Pick-to-Learn with other continual learning approaches could lead to new compelling ways of obtaining self-certified predictors.
Appendix A Table of symbols
We summarize the notation and their definitions in the following table.
Table 2: Summary table of the symbols used in this paper.
Symbol Definition
$\operatorname{\mathcal{X}}$ A feature space
$\operatorname{\mathcal{Y}}$ A label space
$S$ A dataset composed of datapoints in $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$
$\operatorname{\mathcal{D}}$ A distribution over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$
$\operatorname{\mathfrak{D}}$ A meta-distribution over the space of distributions $\operatorname{\mathcal{D}}$
$\Theta$ A set of learnable parameters $\theta$
$f_{\theta}$ A function $f_{\theta}:\operatorname{\mathcal{X}}\to\operatorname{\mathcal{Y}}$ parametrized by $\theta\in\Theta$
AA An learning algorithm that returns parameters $\theta\in\Theta$
$\operatorname{\mathbf{i}}$ The indices of the compression set $S^{\operatorname{\mathbf{i}}}$
$\operatorname{\mathbf{i}}^{{\rm\bf c}}$ The indices of the compression set’s complement $S^{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$
$\operatorname{|\mathbf{i}|}$ The size of the compression set
$\operatorname{\mathbf{j}}$ The indices of the second compression set $S^{\operatorname{\mathbf{j}}}$ used by CoP2L
$\operatorname{\mathscr{P}}(n)$ The powerset of $\{1,\ldots,n\}$
$\mu$ A message
$\Sigma$ The alphabet of symbols $\sigma$ used to create the message $\mu$
$\Sigma^{*}$ The set of all possible sequences of symbols $\sigma\in\Sigma$
$\mathscr{M}(\operatorname{\mathbf{i}})$ The set of possible messages in $\Sigma^{*}$ given a sequence $\operatorname{\mathbf{i}}$
$\operatorname{\mathscr{C}}$ The compression function
@@ Appendix A Table of symbols
$\Psi$ The sample-compression bound of Equation˜5
$\ell$ A loss function $\ell:\Theta\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\to[0,1]$
$\operatorname{\widehat{\mathcal{L}}}_{S}$ The empirical loss over a dataset $S$ given a loss function $\ell$
$\operatorname{\widehat{\mathcal{L}}}_{S}^{\ \operatorname{\mathbf{i}}}$ The empirical loss over a dataset $S^{\operatorname{\mathbf{i}}}$ given a loss function $\ell$
$\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}$ The true loss over a distribution $\operatorname{\mathcal{D}}$ given a loss function $\ell$
Appendix B Compression and Reconstruction Functions
In line with the prevalent literature, we introduced the sample compression framework based on a compression function providing a compression set and optionally a message. In the case of CoP2L, the reconstruction scheme relies on two compression sets $(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}})$ and a message pair $(\mu_{1},\mu_{2})$, which we explain later in this section. [^3] The idea of using multiple compression sets appeared in Marchand and Shawe-Taylor [2002] and Marchand et al. [2003], but in a setting without a message.